OpenStudy (anonymous):

I'm studying for the sat, and heres a question I'm confused about. If the average of A, B, and C is P, what is C in terms of B,C, and P? Could someone walk me through this?

OpenStudy (anonymous):

I have this equation.. A+B+C/3 = P, I just dont know how to proceed.

OpenStudy (raden):

(A+B+C)/3 = P on both sides multiply by 3, we get A+B+C = 3P or C+(A+B) = 3P subtracted by (A+B) on both sides C = 3P - (A+B)

OpenStudy (anonymous):